Circular No. 8106

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003cy AND 2003cz Further to IAUC 8105, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and S. Brar report the discovery of two apparent supernovae:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cy Mar. 30.14 10^{h}46^{m}04.^{s}27 +27^{o}20′57′′.2 19.0 4′′.4 E, 3′′.3 S
2003cz Apr. 2.14 12 05 53.57 +33 27 13.0 20.2 6′′.8 E, 1′′.2 N

Additional magnitudes: SN 2003cy, Mar. 22.14, [21.1; Apr. 2.14, 18.5. SN 2003cz, 2002 Apr. 9.14, [21.4; 20.14, [21.0; 2003 Mar. 21.46, 19.8.

SUPERNOVAE 2003bv AND 2003cq Further to IAUC 8105, T. Matheson, P. Challis, and R. Kirshner re-

port that a spectrum of SN 2003cq (cf. IAUC 8103), obtained by M. Calkins on Apr. 3.28 UT with the 1.5-m telescope, shows it to be a type-Ia supernova about a week before maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9978 km/s for the host galaxy, the supernova expansion velocity is 13200 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 7 ± 2 days before maximum light. A spectrum (range 370-750 nm) of SN 2003bv (cf. IAUC 8093), obtained by L. Macri on Mar. 25.17, shows it to be a type-II supernova. The spectrum consists of a blue continuum, with broad hydrogen emission lines. The absorption component of a P-Cyg profile is not evident, but cannot be excluded due to the poor signal-to-noise ratio of the spectrum.

COMETS C/2002 W6 AND C/2002 W7 (SOHO) Further to IAUC 8103, D. Hammer provides measurements for two

additional Kreutz sungrazing comets found on SOHO C2 website images by R. Matson (C/2002 W6) and D. Evans (C/2002 W7). C/2002 W7 was also visible in the C3 coronagraph. Reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2003-G07.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 W6 Nov. 20.351  15^{h}40.^{m}9  −21^{o}17′
C/2002 W7  21.821  15 49.0  −22 29
2003 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green